OpenStudy (anonymous):

3. What is a polynomial function in standard form with zeroes 1,2,-3, and -1? the choices are g(x)=x^4 +x^3-7x^2-x+6, x^4 +x^3+7x^2-x+6,x^4 +x^3-7x^2-x-6,and x^4 -x^3-7x^2-x+6

hero (hero):

If a polynomial has the given zeroes, then x = 1 x = 2 x = -3 x = -1

hero (hero):

Using algebra, we can re-write the above to say that x - 1 = 0 x - 2 = 0 x + 3 = 0 x + 1 = 0 Do you understand how I got that?

OpenStudy (anonymous):

yea i understand you moved the number over with x

hero (hero):

Because after that you'll realize that the expressions on the left are the factors of the function of interest, and by way of the zero product property we can multiply them together 0 = (x - 1)(x - 2)(x + 3)(x + 1)

OpenStudy (anonymous):

so you substitute the zeros for x?

hero (hero):

No, you multiply the factors on the right completely.

hero (hero):

The factors equal zero when the function equals zero. We don't know yet if the function is g(x)

hero (hero):

If it were g(x), then we could set g(x) = (x - 1)(x - 2)(x + 3)(x + 1)

hero (hero):

But we have to multiply the right side first before deciding which expression is correct.

OpenStudy (anonymous):

how do you do that?

hero (hero):

That depends on how your teacher wants you to do it. If you're doing problems of this type, you are probably expected by your teacher to already understand what it means to multiply factors.

hero (hero):

To do multiply them manually, you would just continue to apply multiplication rules for binomials and polymials until you have a completely expanded expression.

hero (hero):

And if you were doing it manually, then I believe most people would first multiply (x - 1)(x - 2) then (x + 3)(x + 1)

hero (hero):

And then afterwards multiply the products of both together.

hero (hero):

@domanboy97 I assume you are multiplying them together?

OpenStudy (anonymous):

yes

hero (hero):

Let me know what you get

OpenStudy (anonymous):

i got my dad to help me and we got x^4 -x^3-7x^2-x+6 is this right?

hero (hero):

It's not correct though.

hero (hero):

Show me the work you and your dad did to multiply everything out.

hero (hero):

We can figure out the mistakes together.

OpenStudy (anonymous):

i just did it again and got this x^4 +x^3-7x^2-x+6

hero (hero):

I would still like to see the work you did on it.

hero (hero):

I would like to see how you multiplied it all together. There are different techniques you can use to do it. I'm trying to see which of them you used.

hero (hero):